# Interruption Handling for Conversational Robots

### v1

@@ Abstract
Interruptions, a fundamental component of human communication, can enhance the dynamism and effectiveness of conversations, but only when effectively managed by all parties involved. Despite advancements in robotic systems, state-of-the-art systems still have limited capabilities in handling user-initiated interruptions in real-time. Prior research has primarily focused on post hoc analysis of interruptions. To address this gap, we present a system that detects user-initiated interruptions and manages them in real-time based on the interrupter’s intent (i.e., cooperative agreement, cooperative assistance, cooperative clarification, or disruptive interruption). The system was designed based on interaction patterns identified from human-human interaction data. We integrated our system into an LLM-powered social robot and validated its effectiveness through a timed decision-making task and a contentious discussion task with 21 participants. Our system successfully handled 93.69% (n=104/111) of user-initiated interruptions. We discuss our learnings and their implications for designing interruption-handling behaviors in conversational robots.
Keywords: human-robot interaction, conversational robots
@@ I Introduction
Interruptions—occurring when a listener attempts to take the floor before the speaker’s utterance is complete—are a fundamental aspect of human communication. [2]. Interruptions can be used to signal understanding (cooperative agreement), assist the speaker (cooperative assistance), seek clarification when needed (cooperative clarification), or express disagreement and shift the topic (disruptive) [3], leading to more fluid and fast-paced conversations when effectively used and handled. However, interruptions can disrupt the flow of the conversation, potentially leading to conversation breakdowns and the interrupter feeling excluded if addressed inadequately. Thus, it is critical for conversational robots to be able to detect and manage interruptions on-the-fly.
@@ II-A Interruptions in Human-Human Interaction
Interruptions occur when a listener tries to take the floor before the speaker’s utterance is complete [2]. They are a common aspect of human conversations and can occur more than once per minute in dyadic [9] and group conversations [10]. While most interruptions involve overlapping speech, brief overlapping speech that does not purloin the speaker’s floor—such as during speaker exchanges [11] or verbal backchannels (conversation continuers signaling attention, understanding, and agreement, such as ”uh-huh” or ”yeah”) [12]—is not considered an interruption. Interruptions can be categorized as cooperative or disruptive based on the intention of the interrupter [13]. Cooperative interruptions are intended to aid the completion of the current turn by expressing concurrence or understanding (cooperative agreement), providing a word or idea that the speaker may need (cooperative assistance), or asking the speaker to clarify or elaborate on previous information (cooperative clarification) [14]. Disruptive interruptions occur when the listener challenges the speaker’s control and disrupts the conversational flow to express an opposing opinion (disagreement), further develop the current topic (floor taking), change the subject (topic change), or summarize the speaker’s point to end the turn and avoid unwanted information (tangentialization) [14]. Prior work found that the type and frequency of interruptions are shaped by the conversational and power dynamics of interaction [15]. However, limited work investigated how humans handle interruptions in conversations. Hence, in this work, we analyzed human-human conversations with varying power dynamics between speakers to understand how people handle different types of interruptions.
@@ II-B Interruptions in Human-Agent Interaction
Existing conversational robots have very limited capabilities in handling interruptions. Many systems ignore overlapping speech altogether (e.g., [16]), while others treat any overlapping speech as interruption (e.g., [17]) or rely on designed explicit turn-taking signaling (e.g., user touches the robot’s head [4]) to detect user-initiated interruptions. Popular commercial voice assistants take a step further by relying on wakewords (e.g., “Alexa”) to detect user-initiated interruptions [18]. However, these approaches remain error-prone due to their rigidity (may mistake ambient microphone noise as interruption [2] and rely on users to remember the wakeword [19]), and are unnatural in human conversations. Recent advances, such as Alexa’s follow-up mode and Siri’s ability to handle back-to-back requests, though limited, reflect ongoing efforts to improve the fluidity of the conversation. However, these systems still struggle with managing complex, multi-turn interruptions [19], yielding the floor to all interruptions equally without assessing intent. Prior work in human-agent interaction recommended distinguishing between cooperative and disruptive interruptions to help determine how the agent should deal with the interruption [8]. Multi-modal features (i.e., acoustic profiles, head activity, gaze behavior, lexical features, and facial expressions) have been used to automatically classify interruptions as cooperative or competitive [7]. However, prior efforts for handling these cooperative and disruptive interruptions have been limited to post hoc evaluation of a dataset [2]; to the best of our knowledge, no existing conversational agent have the ability to handle user-initiated interruptions based on the context of the interruption in real-time. In this work, we leverage natural language understanding capabilities of large language models (LLMs) to classify and handle user-initiated interruptions in real-time. By categorizing interruptions into four finer-grained categories—cooperative agreement, assistance, clarification, and disruptive—our system tailors handling strategies to match the context and user intention behind the interruption.
@@ III Interruption Handling in Human Interaction
There is limited prior research on interruption handling in human-robot and human-human interaction, offering little guidance for designing conversational robots. To address this, we analyzed natural human conversations to identify human behavior patterns when handling interruptions.
@@ III-A Human Conversational Data
  • Discussions ($n_{D}=106$ interruptions) represent causal settings with no clear power dynamics. We analyzed two dyadic conversations and three multi-party discussions.
@@ III-B Analysis
We coded the videos for 1) intention of the interrupter—either cooperative (i.e., agreement, assistance, or clarification) or disruptive—and 2) the strategy individuals use to manage interruptions—yield or hold the floor. Interaction patterns were created based on the sequence of these two states. One coder transcribed and coded the interruptions, while a secondary coder independently coded $10\%$ to assess inter-coder reliability (Cohen’s Kappa $=1$).
Figure 1: Interaction patterns of interruption handling in human conversations. The highlighted strategies were implemented in our system (Fig. 2).
@@ III-C Interaction Patterns
We identified following interaction patterns (Fig. 1):
@@ III-C3 Cooperative: clarification $\rightarrow$ hold floor or yield
Figure 2: Interruption handling system diagram illustrating how the handling strategy is determined given overlapping speech.
@@ IV Interruption Handling System for Conversational Robots
1– Code: https://github.com/intuitivecomputing/interruption-handling-system/.
@@ IV-A User-Initiated Interruptions Detection
We detect interruptions by monitoring for simultaneous speech. As overlapping speech towards the end of turn is common during speaker floor exchanges and is not considered an interruption [11]. So, if at the time of the overlapping speech, the robot has less than two seconds of planned speech left, the system simply ignores it and finish-up (Line 4). Verbal backchannels are also not considered interruptions; however, distinguishing between verbal backchannel and short interruption (e.g., “No”) requires identifying the intent of the interrupter, detailed in the next section.
@@ V Evaluation Study: Methodology
We conducted a user study to validate our system by integrating the interruption handling system into an LLM-powered social robot. We implemented basic social robot interaction behavior by hand-crafting a bank of facial expressions (happy, satisfied, excited, interested, surprised, and thinking) and head positions (left gaze, right gaze, look at screen, left nod, right nod, thinking), idle behaviors, gaze aversion, and leveraged an LLM (GPT-4o-2024-05-13) to generate contextualized robot speech and select fitting facial expressions, head movements, and task actions 11footnotemark: 1.
@@ V-A Inducing User-Initiated Interruptions
    Robot persona: We designed the robot to have a feminine name (“Luna”), a feminine voice (Google en-US-Standard-H), gaze aversion while it is talking, and generally positive facial expressions (see Appendix Fig. 1). Prior research highlights that gender significantly affects interruption patterns; women are more likely to be interrupted than men, particularly when they avoid direct eye contact with their conversation partners [14]. Additionally, women tend to smile more during conversations, which has been linked to a higher likelihood of being interrupted [14].
@@ V-B Study Procedure
After providing consent, participants completed a demographic questionnaire and watched a 1-minute sample interaction video. They then engaged in a practice task, designed to familiarize them with the robot, answering four factual questions about space exploration. Following the practice task, participants completed two tasks (in random order) and completed a post-task questionnaire after each task. The study concluded with a semi-structured interview aimed to understand their overall experience working with the robot and to collect feedback on robot interruption handling.
@@ V-D Metrics
We constructed two subjective scales (see Appendix B for details) from the post-task questionnaire to explore how interruption handling might affect the user’s:
@@ V-E Coding
One coder transcribed all interactions and extracted conversation snippets with overlapping speech. For ground truth and fine-grained understanding of interruptions, two coders independently analyzed 100% of the interruptions, classifying them by type (disruptive, cooperative agreement, assistance, or clarification) and by attributes such as statement (opinion or non-opinion), question (opinion or factual), or verbal back-channel, and assessed whether each interruption was successfully addressed without causing a conversation breakdown. Cohen’s kappa was $0.92$, and coders resolved conflicts through discussion.
Figure 3: Example conversation where the system handles various types of compounded interruptions. Overlap 3 is an example of a robot-initiated interruption. We used the Jeffersonian transcription system [22].
@@ VI-A The majority of interruption attempts were disruptive
Amongst the interruptions handled by the system ($n=111$), participants used wakeword(s) in $47$ interruptions (decision-making: $n=36$, discussion: $n=11$). Our system is designed to immediately yield when a wakeword is “heard” during overlapping speech ($n=43$). It treats all interruption attempts containing wakeword(s) as strong disruptive interruptions without passing the intent classifier module (SR2, Line 4 in Fig. 2). Due to speech recognition errors, wakewords were not “heard” in $4$ out of $47$ cases. Additionally, interruptions made when less than two seconds of robot-planned speech remained ($n=13$) also bypassed the intent classifier module (SR3, Line 4 in Fig. 2). Of the $55$ interruption attempts that passed through the intent classification module222111111111111 total handled interruptions minus $43$ with wakeword(s) detected, and $13$ occurring near the end of robot-planned speech, the module correctly classified the intent of $44$ interruptions ($80.00\%$). Overall, including interruptions with wakeword(s) detected ($n=43$), our system incorrectly classified the intent in $11$ out of $98$ ($11.22\%$) interruptions, meaning $88.78\%$ were correctly classified.
@@
Disruptive wakeword not used Clarification LLM error handling [^4] LLM failed to generate correctly formatted robot behavior, so the robot performed default LLM error handling. See Appendix A.8 for details. 1 $100\%$
@@ VI-D Speech recognition errors caused the majority of interruption handling failures
Nonetheless, based on the post-study interviews, participants generally perceived their experience with the robot positively ($n=15/21$, describing it as “good”, “nice”, or “helpful”). Among those who recalled initiating interruptions ($n=18$), fourteen found it “easy” to interrupt the robot or were generally satisfied with how it handled interruptions, while four found it “hard”—one because the system failed to detect five interruptions due to network error, and three because they wanted the robot to always yield immediately.
@@ VII Discussion
In this work, inspired by interaction patterns observed in human conversations, we designed and implemented an interruption handling system for conversational robots. We validated with users ($n=21$) that our system successfully handled $93.69\%$ of the interruptions. Next, we discuss design implications based on insights from our exploration to help guide future designs of interruption handling behavior in conversational robots.
@@ VII-B Breaking Habits: From Scripted to Natural HRI
VII-C Limitations and Future Work
Our evaluation involved only one-time interactions in the lab setting. Future work should explore longer or multi-session interactions in the wild. Additionally, our system does not account for non-verbal interruptions. We observed seven instances where the users’ body language (e.g., opening their mouths) indicated an intention to interrupt the robot without making any sounds. Future work should explore how non-verbal interruption can be leveraged in interruption handling and how non-verbal interruptions should be handled by conversational robots. In addition, our system focuses on interruption handling in dyadic conversations; future work should explore how robots should handle interruptions in multi-party and team situations.

### v2

@@ Abstract
Interruptions, a fundamental component of human communication, can enhance the dynamics and effectiveness of conversations, but only when effectively managed by all parties involved. Despite advancements in robotic systems, state-of-the-art systems still have limited capabilities in handling user-initiated interruptions in real-time. Prior research has primarily focused on post hoc analysis of interruptions. To address this gap, we present a system that detects user-initiated interruptions and manages them in real-time based on the interrupter’s intent (i.e., cooperative agreement, cooperative assistance, cooperative clarification, or disruptive interruption). The system was designed based on interaction patterns identified from human-human interaction data. We integrated our system into an LLM-powered social robot and validated its effectiveness through a timed decision-making task and a contentious discussion task with 21 participants. Our system successfully handled 93.69% (n=104/111) of user-initiated interruptions. We discuss our learnings and their implications for designing interruption-handling behaviors in conversational robots.
@@ I Introduction
Interruptions—occurring when a listener attempts to take the floor before the speaker’s utterance is complete—are a fundamental aspect of human communication [8]. Interruptions can be used to signal understanding (cooperative agreement), assist the speaker (cooperative assistance), seek clarification when needed (cooperative clarification), or express disagreement and shift the topic (disruptive) [23], leading to more fluid and fast-paced conversations when effectively used and handled. However, interruptions can disrupt the flow of the conversation, potentially leading to conversation breakdowns and the interrupter feeling excluded if addressed inadequately. Thus, it is critical for conversational robots to be able to detect and manage interruptions on-the-fly [23].
@@ II-A Interruptions in Human-Human Interaction
Interruptions occur when a listener tries to take the floor before the speaker’s utterance is complete [8]. They are a common aspect of human conversations and can occur more than once per minute in dyadic [21] and group conversations [14]. While most interruptions involve overlapping speech, brief overlapping speech that does not purloin the speaker’s floor—such as during speaker exchanges [10] or verbal backchannels (conversation continuers signaling attention, understanding, and agreement, such as “uh-huh” or “yeah”) [18]—is not considered an interruption. Interruptions can be categorized as cooperative or disruptive based on the intention of the interrupter [9]. Cooperative interruptions are intended to aid the completion of the current turn by expressing concurrence or understanding (cooperative agreement), providing a word or idea that the speaker may need (cooperative assistance), or asking the speaker to clarify or elaborate on previous information (cooperative clarification) [13]. Disruptive interruptions occur when the listener challenges the speaker’s control and disrupts the conversational flow to express an opposing opinion (disagreement), further develop the current topic (floor taking), change the subject (topic change), or summarize the speaker’s point to end the turn and avoid unwanted information (tangentialization) [13]. Prior work found that the type and frequency of interruptions are shaped by the conversational and power dynamics of interaction [3]. However, limited work investigated how humans handle interruptions in conversations. Hence, in this work, we analyzed human-human conversations with varying power dynamics between speakers to understand how people handle different types of interruptions.
@@ II-B Interruptions in Human-Agent Interaction
Existing conversational robots have very limited capabilities in handling interruptions. Many systems ignore overlapping speech altogether (e.g., [1]), while others treat any overlapping speech as interruption (e.g., [4]) or rely on designed explicit turn-taking signaling (e.g., user touches the robot’s head [6]) to detect user-initiated interruptions. Popular commercial voice assistants take a step further by relying on wakewords (e.g., “Alexa”) to detect user-initiated interruptions [16]. However, these approaches remain error-prone due to their rigidity (may mistake ambient microphone noise as interruption [8] and rely on users to remember the wakeword [15]), and are unnatural in human conversations. Recent advances, such as Alexa’s follow-up mode and Siri’s ability to handle back-to-back requests, though limited, reflect ongoing efforts to improve the fluidity of the conversation. However, these systems still struggle with managing complex, multi-turn interruptions [15], yielding the floor to all interruptions equally without assessing intent. Prior work in human-agent interaction recommended distinguishing between cooperative and disruptive interruptions to help determine how the agent should deal with the interruption [19]. Multi-modal features (i.e., acoustic profiles, head activity, gaze behavior, lexical features, and facial expressions) have been used to automatically classify interruptions as cooperative or competitive [24]. However, prior efforts for handling these cooperative and disruptive interruptions have been limited to post hoc evaluation of a dataset [8]; to the best of our knowledge, no existing conversational agent has the ability to handle user-initiated interruptions based on the context of the interruption in real-time. In this work, we leverage natural language understanding capabilities of large language models (LLMs) to classify and handle user-initiated interruptions in real-time. By categorizing interruptions into four finer-grained categories—cooperative agreement, assistance, clarification, and disruptive—our system tailors handling strategies to match the context and user intention behind the interruption.
@@ III Interruption Handling in Human Interaction
In HRI, robot behaviors are often designed based on human behavior patterns observed in human-human interaction (e.g., [2, 11]), as this can make robot behavior feel more natural and align better with user expectations. However, there is limited prior research on interruption handling in both human-robot and human-human interaction, providing little guidance for designing conversational robots. To address this, we analyzed natural human conversations to identify human behavior patterns when handling interruptions.
@@ III-A Human Conversational Data
  • Discussions ($n_{D}=106$ interruptions) represent casual settings with no clear power dynamics. We analyzed two dyadic conversations and three multi-party discussions.
@@ III-B Analysis
Informed by the taxonomy of interruption types from prior work and research highlighting the need to tailor interruption handling to the intention (type) of the interruption (detailed in Section II-A), we coded the videos for 1) intention of the interrupter—either cooperative (i.e., agreement, assistance, or clarification) or disruptive—and 2) the strategy individuals use to manage interruptions—yield or hold the floor. Our goal was to identify interaction patterns that reveal how humans handle different types of interruptions. Interaction patterns were created based on the sequence of these two states. One coder transcribed and coded the interruptions, while a secondary coder independently coded $10\%$ to assess inter-coder reliability (Cohen’s Kappa $=1$).
Figure 1: Interaction patterns of interruption handling in human conversations. The highlighted strategies were implemented in our system (Fig. 2). “Ack” denotes acknowledge and “Cont.” denotes continue.
@@ III-C Interaction Patterns
We identified the following interaction patterns (Fig. 1):
@@ III-C3 Cooperative: clarification $\rightarrow$ hold floor or yield
Figure 2: Metro-map-inspired diagram of the interruption handling system, illustrating how user speech flows through the interruption detection, intent classification, and interruption handling modules. It demonstrates how the system selects the handling strategy based on the predicted user intention given overlapping speech between the user and the robot. We use “Line $<$color$>$” to refer to different interruption handling paths in the figure. “Ack” denotes acknowledge and “Cont.” denotes continue.
@@ IV Interruption Handling System for Conversational Robots
1– Code: https://github.com/intuitivecomputing/interruption-handling-system.
@@ IV-A User-Initiated Interruptions Detection
We detect interruptions by monitoring for simultaneous speech. As overlapping speech towards the end of turn is common during speaker floor exchanges and is not considered an interruption [10]. So, if at the time of the overlapping speech, the robot has less than two seconds of planned speech left, the system simply ignores it and “finish-up” (Line 4). Verbal backchannels are also not considered interruptions; however, distinguishing between verbal backchannel and short interruption (e.g., “No”) requires identifying the intent of the interrupter, detailed in the next section.
@@ V Evaluation Study: Methodology
We conducted a user study to validate our system by integrating the interruption handling system into an LLM-powered social robot. We implemented basic social robot interaction behavior by hand-crafting a bank of facial expressions (happy, satisfied, excited, interested, surprised, and thinking) and head positions (left gaze, right gaze, look at screen, left nod, right nod, thinking), idle behaviors, gaze aversion, and leveraged an LLM (GPT-4o-2024-05-13) to generate contextualized robot speech and select fitting facial expressions, head movements, and task actions (additional details provided in supplemental materials\@footnotemark Section II).
@@ V-A Inducing User-Initiated Interruptions
    Robot persona: We designed the robot to have a feminine name (“Luna”), a feminine voice (Google en-US-Standard-H), gaze aversion while it is talking, and generally positive facial expressions (see supplemental materials\@footnotemark Fig. 1). Prior research highlights that gender significantly affects interruption patterns; women are more likely to be interrupted than men, particularly when they avoid direct eye contact with their conversation partners [13]. Additionally, women tend to smile more during conversations, which has been linked to a higher likelihood of being interrupted [13].
@@ V-B Study Procedure
We set up two cameras to record the participants’ interactions with the robot from both the front-view and back-view during the study (see Fig. 3). After providing consent, participants completed a demographic questionnaire and watched a 1-minute sample interaction video. They then engaged in a practice task, designed to familiarize them with the robot, answering four factual questions about space exploration. Following the practice task, participants completed two tasks (in random order) and completed a post-task questionnaire after each task. The study concluded with a semi-structured interview aimed to understand their overall experience working with the robot and to collect feedback on robot interruption handling.
@@ V-D Metrics
We constructed two subjective scales (see supplemental materials\@footnotemark Section III for details) from the post-task questionnaire to explore how interruption handling might affect the user’s:
@@ V-E Coding
One coder transcribed all interactions and extracted conversation snippets with overlapping speech. For ground truth and fine-grained understanding of interruptions, two coders independently analyzed 100% of the interruptions, classifying them by type (disruptive, cooperative agreement, assistance, or clarification) and by attributes such as statement (opinion or non-opinion), question (opinion or factual), or verbal back-channel, and assessed whether each interruption was successfully addressed, meaning that the robot effectively responded to the interrupter without causing a conversation breakdown. Cohen’s kappa was $0.92$, and coders resolved conflicts through discussion.
Figure 3: Example conversation where the system handles various types of compounded interruptions. Overlap 3 is an example of a robot-initiated interruption. We used the Jeffersonian transcription system [17]. “R” denotes robot and “P” denotes participant.
@@ VI-A The majority of interruption attempts were disruptive
Amongst the interruptions handled by the system ($n=111$), participants used wakeword(s) in $47$ interruptions (decision-making: $n=36$, discussion: $n=11$). Our system is designed to immediately yield when a wakeword is “heard” during overlapping speech ($n=43$). It treats all interruption attempts containing wakeword(s) as strong disruptive interruptions without passing the intent classifier module (SR2, Line 4 in Fig. 2). Due to speech recognition errors, wakewords were not “heard” in $4$ out of $47$ cases. Additionally, interruptions made when less than two seconds of robot-planned speech remained ($n=13$) also bypassed the intent classifier module (SR3, Line 4 in Fig. 2). Of the $55$ interruption attempts that passed through the intent classification module222111111111111 total handled interruptions minus $43$ with wakeword(s) detected, and $13$ occurring near the end of robot-planned speech., the module correctly classified the intent of $44$ interruptions ($80.00\%$). Overall, including interruptions with wakeword(s) detected ($n=43$), our system incorrectly classified the intent in $11$ out of $98$ ($11.22\%$) interruptions, meaning $88.78\%$ were correctly classified.
@@
Disruptive wakeword not used Clarification LLM error handling [^4] LLM failed to generate correctly formatted robot behavior, so the robot performed default LLM error handling. See supplemental materials\@footnotemark Section II for details. 1 $100\%$
@@ VI-D Speech recognition errors caused the majority of interruption handling failures
Nonetheless, based on the post-study interviews, participants generally perceived their experience with the robot positively ($n=15/21$, describing it as “good”, “nice”, or “helpful”). Among those who recalled initiating interruptions ($n=18$), 14 found it “easy” to interrupt the robot or were generally satisfied with how it handled interruptions, while four found it “hard”—one because the system failed to detect five interruptions due to network error, and three because they wanted the robot to always yield immediately.
@@ VII Discussion
In this work, inspired by interaction patterns observed in human conversations, we designed and implemented an interruption handling system for conversational robots. To the best of our knowledge, no existing robotic system has integrated intention classification into its interruption handling framework. Most existing conversational robots either ignore interruptions or always yield to any overlapping speech. Systems that ignore interruptions are prone to conversational breakdowns as they disregard the user’s intention to shift the conversation. Alternatively, systems that always yield break the interaction flow when they yield to non-disruptive overlapping speech (e.g., cooperative agreement, cooperative assistance, self-talk, and backchannels), which can result in user frustration over time. The distribution of types of interruptions depends on the task context (i.e., a higher number of disruptive interruptions in our survival task, likely due to the time constraint). In our study, a system that ignores interruptions would fail to handle cooperative clarification (n = 8) and disruptive interruptions (n = 73), leading it to successfully handle only 27.03% of the interruptions. A system that always yields would fail to handle cooperative agreement (n = 16) and cooperative assistance (n = 1), causing it to successfully handle 84.68% of the interruptions. In comparison, our proposed intention-based system successfully handled 93.69% of the interruptions. This shows the benefit of intention-based interruption handling. Next, we discuss design implications based on insights from our exploration to help guide future designs of interruption handling behavior in conversational robots.
@@ VII-B Breaking Habits: From Scripted to Natural HRI
VII-C Limitations
We designed our interruption handling framework based on observations from human-human conversation data from YouTube. However, the sample size was small and limited to a few specific settings. Additionally, our HRI study focused solely on robot as cognitive aid in timed decision-making and contentious discussion scenarios. Future work is needed to explore interruption handling in physical tasks, as well as how interruption handling can be personalized and tailored based on user behavior (e.g., repeated interruptions), user preferences (e.g., their perceived role of the robot), and task context (e.g., time constraint) for designing more comprehensive interruption management for HRI. Furthermore, our study used convenient sampling and did not design interruption handling to support special population. Future work should explore how interruption handling may be designed to improve interactions for special populations (e.g., older adults with neurodegenerative disease). Finally, our evaluation was limited to one-time interactions in the lab setting. Future work should explore longer-term, multi-session interactions in the wild.
Our system also had a few limitations. As shown in our study, speech recognition errors were the primary cause of failure for the intention classification module. While the LLM absorbed some speech recognition errors (some errors did not affect the LLM’s ability to understand context), our current system cannot detect discrepancies between the transcribed speech and actual user speech. Future work could investigate using multimodal inputs with multimodal LLMs as an additional information source. Additionally, our system does not account for non-verbal interruptions. We observed seven instances where the users’ body language (e.g., opening their mouths) indicated an intention to interrupt the robot without making any sounds. Future work should explore how non-verbal interruption can be leveraged in interruption handling and how non-verbal interruptions should be handled by conversational robots. In addition, our system focuses on interruption handling in dyadic conversations; future work should explore how robots should handle interruptions in multi-party and team situations.
Acknowledgments
This work was supported by the National Science Foundation award #2141335.
AI Statement. This paper has been proofread by a language model. The authors verified that the resulting content accurately reflects their original intent.
CRediT author statement. Shiye Cao: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing, Visualization, Project Administration. Jiwon Moon: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing, Visualization. Amama Mahmood: Conceptualization, Methodology, Formal analysis, Investigation, Data curation, Writing, Visualization. Victor Nikhil Antony: Methodology, Software, Writing. Ziang Xiao: Conceptualization, Methodology, Writing-Review and editing. Anqi Liu: Conceptualization, Methodology, Writing-Review and editing. Chien-Ming Huang: Conceptualization, Methodology, Writing-Review and editing, Visualization, Supervision, Funding acquisition.